OpenStudy (anonymous):

vectors : if a= i+2j and b=3i-5j find the magnitude of a.

OpenStudy (anonymous):

\[\left| \left| a \right| \right| = \sqrt{1^{2}+ 2^{2}}\]

OpenStudy (anonymous):

how did you get rid of the i and j ?

OpenStudy (anonymous):

@Kanwar245

OpenStudy (anonymous):

I used the formula of magnitude

OpenStudy (anonymous):

here I used i and j components which is this case are 1 and 2 respectively

OpenStudy (anonymous):

ok thankss